# Optimal Consumption and Portfolio Choice with No-Borrowing Constraint in the Kim-Omberg Model

### 03 Mar 2026 v1

@@ 2.2. From a Dynamic to a Static Budget Constraint
$\pi_{t}=\frac{1}{\sigma\mathcal{H}_{t}}\left(\tilde{Z}_{t}+\frac{\beta_{t}}{\sigma}(\tilde{X}_{t}+\tilde{K}_{t})\right).$
As $(\mathcal{H}_{t})_{t}$, $(\beta_{t})_{t}$, $(\tilde{X}_{t})_{t}$ and $(\tilde{K}_{t})_{t}$ are continuous, these processes are $\mathbb{P}$-a.s. locally bounded. Furthermore, $\int_{0}^{T}\tilde{Z}_{t}^{2}dt<\infty$ $\mathbb{P}$-a.s. for all $T>0$, by definition of the $L^{1}$-solution to (2.11). Thus, $\int_{0}^{T}\pi_{t}^{2}dt<\infty$ $\mathbb{P}$-a.s. for all $T>0$, which implies $(\pi,c)\in\mathcal{A}(x)$. ∎
@@ 3.1. Derivation of the Dual Problem
(3.9) $d\beta_{t}=\kappa(\overline{\beta}-\beta_{t})\,dt+\sigma_{\beta}\,dW_{t}^{\beta},\quad t>0,\quad\beta_{0}=\beta\in\mathbb{R},$
@@ 3.2. Derivation of the Auxiliary Optimal Stopping Problem
where $\tilde{u}^{\prime}(z)=-z^{1/\gamma}$, and where we take the infimum over $\bar{\mathbb{F}}$-stopping times $\tau\geq 0$, $Z^{1}$ evolves as in (3.12), and $\beta$ as in (3.9). We expect $v$ to be such that $\tilde{V}_{z}=v$ on $\mathcal{O}$. Theorem 4.1 below will indeed prove that such a relation holds true and that an optimizer for $v$ is in one-to-one correspondence to an optimizer for $\tilde{V}$. In the following, we shall study (3.13) and characterize its optimal policy. In order to achieve this, it is convenient to perform a change of measure to remove the martingale $(M_{t})_{t}$ (cf. (2.3)) from the stopping functional. This leads to the next proposition whose proof is postponed to the appendix.
@@
$d\hat{\beta}_{t}$ $=\sigma_{\beta}\,dW^{\beta,\mathbb{Q}}_{t}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})-\rho\,\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)\,dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
@@
(3.17) $d\hat{\beta}_{t}$ $=\sigma_{\beta}dW_{t}^{\beta,\mathbb{Q}}+(\kappa(\overline{\beta}-\hat{\beta}_{t})-\rho\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta})dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
@@ 3.3. Preliminary Properties of the Optimal Stopping Value Function
$\gamma>\max\bigg\{1,\frac{|\rho|\sigma_{\beta}}{\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)}\bigg\}.$
Notice that $\kappa+\frac{\rho\sigma_{\beta}}{\sigma}>0$ due to Assumption 2.1. We then have the following first preliminary finding.
@@
$1<p<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)}{|\rho|\sigma_{\beta}}\bigg\},$
upon noticing that, by Assumption 3.2, we have $\gamma\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)>|\rho|\sigma_{\beta}$. Therefore, by Vitali’s Convergence Theorem and continuity of $\beta\mapsto\hat{Z}_{t}^{z,\beta}$, we conclude that
@@ 3.4. Optimal Stopping Boundary and Regularity of the Value Function
where $a:=\kappa+\rho\frac{\sigma_{\beta}}{\sigma}>0$ by Assumption 2.1.
@@
(3.36) $(\mathcal{L}v)(z,\beta)=\frac{1}{2}\frac{\beta^{2}}{\sigma^{2}}(2zv_{z}+z^{2}v_{zz})+\frac{1}{2}\sigma_{\beta}^{2}v_{\beta\beta}+(\delta-r)zv_{z}+\kappa(\overline{\beta}-\beta)v_{\beta}-\rho\frac{\beta}{\sigma}\sigma_{\beta}\,(v_{\beta}+zv_{z\beta}),$
@@
Given that $Corr(W_{t}^{\mathbb{Q}},W_{t}^{\beta,\mathbb{Q}})=\rho t$, $\rho\in[-1,1]$, we can write
$W_{t}^{\beta,\mathbb{Q}}=\rho\,W_{t}^{\mathbb{Q}}+\sqrt{1-\rho^{2}}\,\widehat{W}_{t}^{2,\perp},\quad t\geq 0,$
where $\widehat{W}^{2,\perp}$ is a standard Brownian motion independent of $W^{\mathbb{Q}}$. Then (cf. (3.16) and (3.17)),
@@
$d\hat{\beta}_{t}=\sigma_{\beta}\rho\,dW_{t}^{\mathbb{Q}}+\sigma_{\beta}\sqrt{1-\rho^{2}}\,d\widehat{W}_{t}^{2,\perp}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})-\rho\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
Notice that (3.36) is not uniformly elliptic. As a matter of fact, denoting by $\Sigma(z,\beta)$ the diffusion matrix associated to (3.16)-(3.17) one has $\det(\Sigma\Sigma^{T}(z,\beta))=\frac{\beta^{2}z^{2}\sigma_{\beta}^{2}}{\sigma^{2}}(1-\rho^{2})$, which vanishes for $\beta=0$. We therefore now check that the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ satisfies the so-called Hörmander’s condition (see, e.g., condition (H) in Section 2.3.2 in [34]). This implies that the second-order infinitesimal generator $\mathcal{L}$ of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ as in (3.36) is hypoelliptic and therefore that $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ is a strong Feller process (see Proposition 4 in [13]).
For $(z,\beta)\in\mathcal{O}$, arbitrary but fixed, we then define the Stratonovich-corrected drift vector field and the diffusion vector fields driven by $W^{\mathbb{Q}}$ and $\widehat{W}^{2,\perp}$ as follows:
$V^{(0)}(z,\beta)=\begin{pmatrix}\dfrac{z\big(\beta^{2}+\rho\sigma\sigma_{\beta}+2\sigma^{2}(\delta-r)\big)}{2\sigma^{2}}\\[8.0pt] \dfrac{-\beta\rho\sigma_{\beta}+\kappa\sigma(\overline{\beta}-\beta)}{\sigma}\end{pmatrix},\qquad V^{(1)}(z,\beta)=\begin{pmatrix}-\dfrac{\beta}{\sigma}z\\[4.0pt] \sigma_{\beta}\rho\end{pmatrix}$,
$V^{(2)}(z,\beta)=\begin{pmatrix}0\\[4.0pt] \sigma_{\beta}\sqrt{1-\rho^{2}}\end{pmatrix}$.
First, consider the case $|\rho|<1$. We denote by $DV^{(i)}$ the Jacobian matrix of the vector $V^{(i)}$, $i=0,1,2$. The Lie-bracket between $V^{(1)}$ and $V^{(2)}$, denoted by $V^{(1)},V^{(2)}]$, is such that
$V^{(1)},V^{(2)}](z,\beta):=-DV^{(1)}V^{(2)}(z,\beta)=-\begin{pmatrix}-\frac{\beta}{\sigma}&-\frac{z}{\sigma}\\[6.0pt] 0&0\end{pmatrix}\begin{pmatrix}0\\[6.0pt] \sigma_{\beta}\sqrt{1-\rho^{2}}\end{pmatrix}=\begin{pmatrix}\dfrac{\sigma_{\beta}\sqrt{1-\rho^{2}}}{\sigma}z\\[6.0pt] 0\end{pmatrix}$.
Since $z>0$, $\sigma_{\beta}>0$, and $|\rho|<1$, the vectors $V^{(2)}(z,\beta)$ and $V^{(1)},V^{(2)}](z,\beta)$ are linearly independent and therefore span $\mathcal{O}$.
Now consider the case $|\rho|=1$. The Lie bracket between $V^{(0)}$ and $V^{(1)}$ is
$V^{(0)},V^{(1)}](z,\beta):=DV^{(1)}V^{(0)}(z,\beta)-DV^{(0)}V^{(1)}(z,\beta)=\begin{pmatrix}-\dfrac{z\kappa(\overline{\beta}-\beta)}{\sigma}\\[6.0pt] \dfrac{\rho\sigma_{\beta}}{\sigma}\big(\kappa\sigma+\rho\sigma_{\beta}\big)\end{pmatrix}$.
@@
$V^{(1)},[V^{(0)},V^{(1)}]](z,\beta)=D[V^{(0)},V^{(1)}]V^{(1)}(z,\beta)-DV^{(1)}[V^{(0)},V^{(1)}]=\begin{pmatrix}\dfrac{\rho\sigma_{\beta}z}{\sigma^{2}}\big(2\kappa\sigma+\rho\sigma_{\beta}\big)\\[6.0pt] 0\end{pmatrix}$,
@@
$\det(V^{(1)},[V^{(1)},[V^{(0)},V^{(1)}]])=-\dfrac{\rho^{2}\sigma_{\beta}^{2}z}{\sigma^{2}}\big(2\kappa\sigma+\rho\sigma_{\beta}\big)<0,$
@@
$2\kappa\sigma+\rho\sigma_{\beta}>(2+\rho)\sigma_{\beta}>0,$
where Assumption 2.1 has been used. Hence, $V^{(1)}$ and $V^{(1)},[V^{(0)},V^{(1)}]]$ are linearly independent and thus span $\mathcal{O}$. Hörmander’s condition is therefore verified and we conclude that the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ is indeed strong Feller. ∎
@@
Since the infinitesimal generator of the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ is hypoelliptic (satisfying Hörmander’s conditions; cf. the proof of Proposition 3.12), the drift and volatilities of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ belong to $C^{\infty}(\mathcal{W})$, and $\tilde{u}^{\prime}(z)+\ell\in C^{\infty}(\mathcal{W})$, Corollary 7 in [36] implies that $v$ is not just a solution to (3.35) in the sense of Schwartz distributions but $v\in C^{\infty}(\mathcal{W})$ and thus solves (3.35) in the classical sense. ∎
@@
By the proof of Lemma 3.12, we know that the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ satisfies the Hörmander’s condition. Given that (modulo a change of measure to remove the quadratic term in $\hat{\beta}$ appearing in the drift of the dynamics for $\hat{Z}$) the coefficients of the evolution of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ are linear and infinitely many times differentiable (see (3.16) and (3.17)), it thus follows from Theorem 9-(iii) and Remark 11 in [3] or Theorem 2.3.3 in [34], among others, that, for any $t>0$, $(\hat{Z}_{t},\beta_{t})_{t}$ admits a transition density that is infinitely many times differentiable in its arguments.
@@
$X_{t}:=e^{-rt}g(\xi_{t})=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{t}^{\infty}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds\,\Big|\,\hat{\mathcal{F}}_{t}\bigg],$
@@
$M_{t}:=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds\,\Big|\,\hat{\mathcal{F}}_{t}\bigg],\quad\text{and}\quad A_{t}:=-\int_{0}^{t}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds.$
Notice that, since $\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rs}|\tilde{u}^{\prime}(\hat{Z}_{s})+\ell|\;ds\bigg]<\infty$ by Proposition 3.3, $(M_{t})_{t}$ is a uniformly integrable $\hat{\mathbb{F}}-$martingale and $dA_{t}=dA^{+}_{t}+dA^{-}_{t}$ with
@@
An immediate consequence of Propositions 3.13, 3.14, (3.36), and the fact that $v=0$ in the interior of $\mathcal{S}$, denoted by $\mathring{\mathcal{S}}$, is the following Corollary.
@@
One has $v\in C^{1}(\mathcal{O})\cap C^{\infty}(\mathcal{W}\cup\mathring{\mathcal{S}})$. Furthermore, $\frac{1}{2}\frac{\beta^{2}}{\sigma^{2}}z^{2}v_{zz}+\frac{1}{2}\sigma_{\beta}^{2}v_{\beta\beta}-\frac{\rho\sigma_{\beta}}{\sigma}\,\beta zv_{z\beta}$ admits a continuous extension to $\overline{\mathcal{W}}$.

### 11 Mar 2026 v3

@@ 2.2. From a Dynamic to a Static Budget Constraint
$d(\mathcal{H}_{t}^{-1}\widehat{X}_{t})=\left(r\mathcal{H}_{t}^{-1}\widehat{X}_{t}+\ell-c_{t}+\frac{\beta_{t}}{\sigma}\left(\mathcal{H}_{t}^{-1}\tilde{Z}_{t}+\frac{\beta_{t}}{\sigma}\mathcal{H}_{t}^{-1}\widehat{X}_{t}\right)\right)dt+\left(\mathcal{H}_{t}^{-1}\tilde{Z}_{t}+\frac{\beta_{t}}{\sigma}\mathcal{H}_{t}^{-1}\widehat{X}_{t}\right)dW_{t},$
or equivalently, upon using (2.21),
$d(\mathcal{H}_{t}^{-1}\widehat{X}_{t})=(r\mathcal{H}_{t}^{-1}\widehat{X}_{t}+\beta_{t}\pi_{t}-c_{t}+\ell)\,dt+\sigma\pi_{t}\,dW_{t},\qquad\mathcal{H}_{0}^{-1}\widehat{X}_{0}=x.$
[…]
(2.23) $X_{t}^{\pi,c}=\mathcal{H}_{t}^{-1}\widehat{X}_{t},\qquad t\geq 0,$
by the uniqueness of the solution to the SDE. Finally, since $\widehat{X}_{t}\geq 0$ and $\mathcal{H}_{t}>0$ for all $t\geq 0$, we have from (2.23)
$X_{t}^{\pi,c}\geq 0\qquad\text{for all }t\geq 0,$
and therefore, $(\pi,c)\in\mathcal{A}(x)$. ∎
@@ 3.1. Derivation of the Dual Problem
(3.9) $d\beta_{t}=\kappa(\overline{\beta}-\beta_{t})\,dt-\sigma_{\beta}\,dW_{t},\quad t>0,\quad\beta_{0}=\beta\in\mathbb{R},$
@@
Remark 3.1.
We emphasize that while the complete-market assumption is needed in order to establish strong duality and recover the solution to the primal problem from the dual formulation, the singular stochastic control problem (3.6) and the associated auxiliary optimal stopping problem introduced and studied below can be formulated and analyzed as in the subsequent sections more generally when the factor process $(\beta_{t})_{t}$ is driven by a Brownian motion $(W_{t}^{\beta})_{t}$ satisfying
$W_{t}^{\beta}=\rho W_{t}+\sqrt{1-\rho^{2}}\,W_{t}^{\perp},\qquad t\geq 0,$
where $(W^{\perp}_{t})_{t}$ is a standard Brownian motion independent of $(W_{t})_{t}$ and $\rho\in[-1,1]$.
@@ 3.2. Derivation of the Auxiliary Optimal Stopping Problem
where $\tilde{u}^{\prime}(z)=-z^{1/\gamma}$, and where we take the infimum over $\mathbb{F}^{W}$-stopping times $\tau\geq 0$, $Z^{1}$ evolves as in (3.12), and $\beta$ as in (3.9). We expect $v$ to be such that $\tilde{V}_{z}=v$ on $\mathcal{O}$. Theorem 4.1 below will indeed prove that such a relation holds true and that an optimizer for $v$ is in one-to-one correspondence to an optimizer for $\tilde{V}$. In the following, we shall study (3.13) and characterize its optimal policy. In order to achieve this, it is convenient to perform a change of measure to remove the martingale $(M_{t})_{t}$ (cf. (2.5)) from the stopping functional. This leads to the next proposition whose proof is postponed to the appendix.
@@
$d\hat{\beta}_{t}$ $=-\sigma_{\beta}\,dW^{\mathbb{Q}}_{t}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})+\,\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)\,dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
@@
(3.17) $d\hat{\beta}_{t}$ $=-\sigma_{\beta}dW_{t}^{\mathbb{Q}}+(\kappa(\overline{\beta}-\hat{\beta}_{t})+\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta})dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
@@ 3.3. Preliminary Properties of the Optimal Stopping Value Function
$\gamma>\max\bigg\{1,\frac{\sigma_{\beta}}{\sigma\bigg(\kappa-\frac{\sigma_{\beta}}{\sigma}\bigg)}\bigg\}.$
Notice that $\kappa-\frac{\sigma_{\beta}}{\sigma}>0$ due to Assumption 2.1. We then have the following first preliminary finding.
@@
$1<p<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa-\frac{\sigma_{\beta}}{\sigma}\bigg)}{\sigma_{\beta}}\bigg\},$
upon noticing that, by Assumption 3.3, we have $\gamma\sigma\bigg(\kappa-\frac{\sigma_{\beta}}{\sigma}\bigg)>\sigma_{\beta}$. Therefore, by Vitali’s Convergence Theorem and continuity of $\beta\mapsto\hat{Z}_{t}^{z,\beta}$, we conclude that
@@ 3.4. Optimal Stopping Boundary and Regularity of the Value Function
where $a:=\kappa-\frac{\sigma_{\beta}}{\sigma}>0$ by Assumption 2.1.
@@
(3.36) $(\mathcal{L}v)(z,\beta)=\frac{1}{2}\frac{\beta^{2}}{\sigma^{2}}(2zv_{z}+z^{2}v_{zz})+\frac{1}{2}\sigma_{\beta}^{2}v_{\beta\beta}+(\delta-r)zv_{z}+\kappa(\overline{\beta}-\beta)v_{\beta}+\frac{\beta}{\sigma}\sigma_{\beta}\,(v_{\beta}+zv_{z\beta}),$
@@
Recall the dynamics of the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ (cf. (3.16) and (3.17)),
@@
$d\hat{\beta}_{t}=-\sigma_{\beta}\,dW_{t}^{\mathbb{Q}}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})+\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
Notice that (3.36) is not uniformly elliptic. As a matter of fact, denoting by $\Sigma(z,\beta)$ the diffusion matrix associated to (3.16)-(3.17) one has $\det(\Sigma\Sigma^{T}(z,\beta))=0$. We therefore now check that the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ satisfies the so-called Hörmander’s condition (see, e.g., condition (H) in Section 2.3.2 in [34]). This implies that the second-order infinitesimal generator $\mathcal{L}$ of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ as in (3.36) is hypoelliptic and therefore that $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ is a strong Feller process (see Proposition 4 in [14]).
For $(z,\beta)\in\mathcal{O}$, arbitrary but fixed, we then define the Stratonovich-corrected drift vector field and the diffusion vector field driven by $W^{\mathbb{Q}}$ as follows:
$V^{(0)}(z,\beta)=\begin{pmatrix}\dfrac{z\big(\beta^{2}-\sigma\sigma_{\beta}+2\sigma^{2}(\delta-r)\big)}{2\sigma^{2}}\\[8.0pt] \dfrac{\beta\sigma_{\beta}+\kappa\sigma(\overline{\beta}-\beta)}{\sigma}\end{pmatrix},\qquad V^{(1)}(z,\beta)=\begin{pmatrix}-\dfrac{\beta}{\sigma}z\\[4.0pt] -\sigma_{\beta}\end{pmatrix}$.
We denote by $DV^{(i)}$ the Jacobian matrix of the vector $V^{(i)}$, $i=0,1$. The Lie bracket between $V^{(0)}$ and $V^{(1)}$, denoted by $V^{(0)},V^{(1)}]$, is such that
$V^{(0)},V^{(1)}](z,\beta):=DV^{(1)}V^{(0)}(z,\beta)-DV^{(0)}V^{(1)}(z,\beta)=\begin{pmatrix}-\dfrac{z\kappa(\overline{\beta}-\beta)}{\sigma}\\[6.0pt] -\dfrac{\sigma_{\beta}}{\sigma}\big(\kappa\sigma-\sigma_{\beta}\big)\end{pmatrix}$.
@@
$V^{(1)},[V^{(0)},V^{(1)}]](z,\beta)=D[V^{(0)},V^{(1)}]V^{(1)}(z,\beta)-DV^{(1)}[V^{(0)},V^{(1)}]=\begin{pmatrix}\dfrac{-\sigma_{\beta}z}{\sigma^{2}}\big(2\kappa\sigma-\sigma_{\beta}\big)\\[6.0pt] 0\end{pmatrix}$,
@@
$\det(V^{(1)},[V^{(1)},[V^{(0)},V^{(1)}]])=-\dfrac{\sigma_{\beta}^{2}z}{\sigma^{2}}\big(2\kappa\sigma-\sigma_{\beta}\big)<0,$
@@
$2\kappa\sigma-\sigma_{\beta}>0,$
in which Assumption 2.1 has been used. Hence, $V^{(1)}$ and $V^{(1)},[V^{(0)},V^{(1)}]]$ are linearly independent and thus span $\mathcal{O}$. Hörmander’s condition is therefore verified and we conclude that the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ is indeed strong Feller. ∎
@@
Since the infinitesimal generator of the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ is hypoelliptic (satisfying Hörmander’s conditions; cf. the proof of Lemma 3.13), the drift and volatilities of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ belong to $C^{\infty}(\mathcal{W})$, and $\tilde{u}^{\prime}(z)+\ell\in C^{\infty}(\mathcal{W})$, Corollary 7 in [37] implies that $v$ is not just a solution to (3.35) in the sense of Schwartz distributions but $v\in C^{\infty}(\mathcal{W})$ and thus solves (3.35) in the classical sense. ∎
@@
By the proof of Lemma 3.13, we know that the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ satisfies the Hörmander’s condition. Given that (modulo a change of measure to remove the quadratic term in $\hat{\beta}$ appearing in the drift of the dynamics for $\hat{Z}$) the coefficients of the evolution of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ are linear and infinitely many times differentiable (see (3.16) and (3.17)), it thus follows from Theorem 9-(iii) and Remark 11 in [3] or Theorem 2.3.3 in [34], among others, that, for any $t>0$, $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ admits a transition density that is infinitely many times differentiable in its arguments.
@@
$X_{t}:=e^{-rt}g(\xi_{t})=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{t}^{\infty}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds\,\Big|\,\mathcal{F}_{t}^{W,\mathbb{Q}}\bigg],$
@@
$M_{t}:=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds\,\Big|\,\mathcal{F}_{t}^{W,\mathbb{Q}}\bigg],\quad\text{and}\quad A_{t}:=-\int_{0}^{t}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds.$
Notice that, since $\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rs}|\tilde{u}^{\prime}(\hat{Z}_{s})+\ell|\;ds\bigg]<\infty$ by Proposition 3.4, $(M_{t})_{t}$ is a uniformly integrable $\mathbb{F}^{W,\mathbb{Q}}$-martingale and $dA_{t}=dA^{+}_{t}+dA^{-}_{t}$ with
@@
An immediate consequence of Propositions 3.14, 3.15, (3.36), and the fact that $v=0$ in the interior of $\mathcal{S}$, denoted by $\mathring{\mathcal{S}}$, is the following Corollary.
@@
One has $v\in C^{1}(\mathcal{O})\cap C^{\infty}(\mathcal{W}\cup\mathring{\mathcal{S}})$. Furthermore, $\frac{1}{2}\frac{\beta^{2}}{\sigma^{2}}z^{2}v_{zz}+\frac{1}{2}\sigma_{\beta}^{2}v_{\beta\beta}+\frac{\sigma_{\beta}}{\sigma}\,\beta zv_{z\beta}$ admits a continuous extension to $\overline{\mathcal{W}}$.